### v16

# Near-Optimal Very Low Rank Approximation at Sublinear Cost

Abstract
Eckart-Young’s theorem implies that even optimal rank-$r$ approximation of an $m\times n$ matrix $M$ for $m\geq n$ is only interesting if there is a gap in the segment of the $r+1$ largest singular values of $M$. If $M$ also has a gap in the segment of the next $p$ largest singular values, then one can compute its rank-$r$ approximation with random sampling and oversampling in nearly linear arithmetic time. For matrices $M$ having such a double gap in their spectra and for $n$ greatly exceeding $(r+p)^{2}$, our algorithm runs in sublinear time, and for a large subclass of the class of such matrices $M$ its output rank-$r$ approximation of $M$ is near-optimal under the spectral norm. We also extend the classical iterative refinement to LRA and analyze and test the resulting algorithm.
Key Words:
low rank approximation, singular values, gaps in the spectrum, sublinear cost algorithms, iterative refinement
2020 Math. Subject Classification:
65F55, 65N75, 65Y20, 68Q25, 68W20
1 LRA of a matrix: the State of the Art and our Novelties
Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra. We first briefly recall the State of the Art, referring to [11, 33, 18] for broader expositions, and then outline our LRA algorithms.
1.1 The State of the Art
1.1.1. LRA and CUR LRA. Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors. [^1] The sketch of a matrix in [36, 33] is its product with a fixed set of vectors. Then we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [33, Sec. 1.3]). Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [35], that is, admit LRA, with which one can operate superfast, aka at sublinear cost – by using much fewer than $mn$ scalars and flops. [^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [24], and we extend this concept to other matrices.
Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r\ll n$ or a fixed target range [^3] Here and hereafter $a\ll b$ means that $b$ greatly exceeds $a$ in context, e.g., $a=o(b)$ as $b\mapsto\infty$. for such a rank (actually, a proper choice of such targets is a challenge [11, Sec. 4.2]) and then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
$X\in\mathbb{C}^{m\times k},Y\in\mathbb{C}^{k\times\ell},~{}Z\in\mathbb{C}^{\ell\times n},$ (1.1)
$k\ll m,~{}\ell\ll n,~{}\min\{k,\ell\}=r,$ (1.2)
and the spectral and/or Frobenius norm of the matrix $M-XYZ$ is small in context, say, close to optimal.
The matrices $X,Y,Z$ have $mk+k\ell+\ell n\ll mn$ entries overall, and we can multiply the matrix $XYZ$ by a vector by using less than $2(mk+k\ell+\ell n)\ll mn$ flops.
If the matrices $X$ and $Z$ are made up of the columns and rows of $M$, that is, are defined by means of pre- and post-multiplication of $M$ by selected coordinate vectors, respectively, then write $X:=C$, $Z:=R$, and $Y:=U$ and call the triplet $\{X,Y,Z\}$ a CUR LRA.
1.1.2. LRA from top SVD, SVD of an LRA, and CUR from SVD. By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, obtain a matrix $M_{r}$ of rank $r$, said to be $r$-truncation of $M$. By virtue of the Eckart-Young-Mirsky theorem [7, Thm. 2.4.8], $M_{r}$ is an optimal rank-$r$ approximation444“Rank- $r$ approximation” means “rank at most $r$ approximation” throughout. of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [7, Fig. 8.6.1]). This is too expensive for computing LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ and then its CUR LRA by using $(m+k)k+(n+\ell)\ell$ scalars, $O(mk^{2}+n\ell^{2})$ flops and $(m+n+r)r$ scalars, $O(mr^{2})$ flops, respectively (see Appendices B and C.2, respectively).
1.1.3. LRA via random sampling and oversampling. Random sampling with oversampling enables fast computation of LRA (see [3, 17, 11, 36, 33, 20], and the references therein). For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [11] successively generates a random test matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
Remark 1.1.
“In practice, the target rank $k$ is rarely known in advance. Randomized algorithms are usually implemented in anadaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [11, Sec. 4.2]).
In the best studied case the test matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables. In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [11, Thm. 10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops. The parameter $p$ and the size of the test matrix $H$ explode as $\epsilon\mapsto 0$, but LRA with $\epsilon$ as large as 1/2 or 1 is still valuable.
In other LRA algorithms (e.g., in [4, 3, 33, 20]) $M$ is both pre- and post-multiplied by random test matrices $F$ and $H$, respectively. In particular, the TYUC algorithm of [33], modifying those of [3, Thm. 4.7] and [3, Thm. 4.3, display 1], and the generalized Nyström approximation GN of [20] first compute the products $FM$ and $MH$ for random test matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$ with the expected Frobenius error norm within a factor of 2 from optimal in the case of random Gaussian test matrices $F$ and $H$ (namely, the norm does not exceed the optimal one for rank-$\frac{k}{2}$ approximation). For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [20]) and $O(\ell mn)$ flops.
The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [11, Secs. 10 and 11], [36, Ch. 6.2], [5, Thm. 3.4]).
1.1.4. Superfast LRA for special subclasses of matrices. For some special but large and important classes of matrices, most notably for Symmetric Positive Semidefinite (SPSD) matrices but also for Distance matrices and a subclass of Cauchy matrices, the superfast randomized algorithms of [19, 1, 12, 14, 2] compute LRA whose expected Frobenius error norm is compared well with the optimal bound. The mere existence of such LRA algorithms is surprising for $n\times n$ SPSD matrices, whose class includes matrices $A^{*}A$, defined by nonsingular $n\times n$ matrices $A$ with $n^{2}$ independent entries. Notice that the superfast algorithm of [15, Part III] for SPSD matrices is deterministic, but its spectral and Frobenius output error norms exceed their optimal values by at most a factor of $n$.
Unfortunately, one cannot compute LRA superfast for every matrix: already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [26, 27, 28]). If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0.
1.2 Our LRA algorithm
Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$. We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$ of size $m\times n$ such that $r<\rho\ll\sqrt{n}$. Our algorithm is superfast and, for a large subclass of $(r,\rho)$-matrices $M$, its output very low rank approximation (VLRA) is near-optimal under spectral norm.
We rely on combining three surprisingly simple observations, namely, that (i) random sampling with oversampling is the only stage where HMT, TYUC, and GN algorithms are not superfast for VLRA; (ii) we can perform this stage superfast by using sparse test matrices (in particular abridged SRHT matrices of Appendix D), and (iii) while the accuracy of the VLRA output in step (ii) tends to be unsatisfactory, we can readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$. Our LRA Alg. 2.1 of Sec. 2 relies on these observations.
We compute at first (1) a crude but reasonably close rank-$\rho$ approximation $M(\rho)$ of $M$ and then (2) the rank-$r$ truncation $M(\rho)_{r}$ of $M(\rho)$, which we output as rank-$r$ approximation of $M$.
LRA via random sampling with oversampling fits this outline, but based on above observations (i)–(iii) we run the algorithm superfast for $\rho\ll\sqrt{n}$ and argue that its output VLRA is near-optimal for a large subclass of $(r,\rho)$-matrices.
We can perform stage (1) of the algorithm superfast by applying any of the HMT, TYUC and GN algorithms for $k:=\rho\ll\sqrt{n}$ with sparse test matrices such that their multiplication by $M$ is superfast (see [13], [33, Sec. 3.9], and our Sec. 3).
Apart from the stage of multiplication by test matrices, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$, and obtain Moore-Penrose pseudoinverses of a $k\times\ell$ matrices ${\bf\Psi Q}$ in [33] and $R_{\epsilon}$ in [20] by using $O(k^{2}\ell)$ flops (see [7, Fig. 8.6.1]). These computations are superfast where $(k+p)^{2}\ll n$ and $k^{2}\ll n$, respectively, for $\ell\leq 2k$, say.
At Stage (2) we apply Alg. B.1. It computes SVD and $r$-truncation of $M(\rho)$ by using about $(m+n)\rho$ scalars and $O(mr^{2})$ flops, that is, $\ll m\sqrt{n}$ scalars and $\ll mn$ flops for $\rho\ll\sqrt{n}$.
By virtue of Thm. 2.3, the spectral norm $||\cdot||:=||\cdot||_{2}$ of the output error matrix of Alg. 2.1 satisfies
$||M-M(\rho)_{r}||\leq||M-M_{r}||+2||M-M(\rho)||,$ (1.3)
that is, Alg. 2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$. For $(r,\rho)$-matrices, this factor can be large only if the superfast computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails. It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse test matrices tend to be unsatisfactory as LRA [13], [33, Sec. 3.9], but in our numerical tests in Sec. 5 accuracy of the output matrices of stage (1) of Alg. 2.1 tended to be raised to near-optimal level at stage (2).
1.3 Iterative LRA
In Sec. 4 we apply our LRA algorithm to the output error matrix, hoping to decrease its norm if it is too large. Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs. The algorithm is essentially reduced to recursive applications of Alg. 2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance. We counter this deficiency by recursively compressing the sums into rank-$r$ matrices, although this limits output accuracy of computed LRAs. The resulting iterative LRA algorithm is a natural extension to LRA of an efficient popular technique of matrix computations; we specify, analyze, and test it in Secs. 4 and 5.
We devised this algorithm to complement and strengthen Alg. 2.1, but it is essentially reduced to recursive applications of Alg. 2.1. Our upper estimates for the accuracy of both algorithms are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
In our tests, the output accuracy of our iterative LRA applied to $(r,\rho)$-matrices improved significantly at the second iteration, little improved afterwards, and overall error norm was never less than with Alg. 2.1 for the upper rank $\rho=4r$.
1.4 Related works
Our study was prompted by extensive work on LRA in the last decades (see [11, 33, 18] and the bibliography therein) and in particular by fast LRA based on random sampling and oversampling. To accelerate these algorithms we use abridged SRHT matrices studied in the papers [26, 27, 28], which in turn extends the earlier study of Gaussian elimination without pivoting in [29, 31].
Our iterative LRA extends the algorithms of [16, 25], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [32, Secs. 3.3.4 and 4.2.5], [10, Ch. 12 and Sec. 20.5], [7, Secs. 3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [32, page 223 and the references on page 225]). Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [22, 30].
Actually, we first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.
1.5 Organization of our paper
We present our accelerated oversampling LRA algorithm in Sec. 2, discuss superfast computation of a crude LRA in Sec. 3, present our iterative LRA algorithm in Sec. 4, and cover our numerical experiments in Sec. 5. We devote our short Sec. 6 to conclusions. In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails. In Appendix B we compute $r$-top SVD of an LRA of a low-rank matrix superfast. In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA. In Appendix D we generate abridged SRHT matrices, whose incorporation as test matrices makes random sampling LRA superfast.
2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
We begin with some background materials.
Hereafter $||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Remark 2.2. In particular we only use Eckart-Young’s (rather than Mirski’s) theorem:
Theorem 2.1.
$\sigma_{r+1}(M)=||M_{r}-M||=\min_{X:~{}\operatorname{rank}(X)\leq r}||M-X||$.
Theorem 2.2.
[7, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
$|\sigma_{j}(M+E)-\sigma_{j}(M)|\leq||E||~{}{\rm for}~{}j=1,\dots,\min\{m,n\}.$
In the following algorithm we elaborate upon the outline of Sec. 1.2 assuming that we are given two black box subroutines that output a crude LRA of a matrix.
Algorithm 2.1.
LRA of $(r,\rho)$-matrices.
Input:
An $m\times n$ matrix $M$, for $n\leq m$, given implicitly by an oracle (black box subroutine) for its multiplication by a vector (cf. Sec. 1.1.1), a target rank $r\ll n$, and a tolerance $\mu$ to the output error norm.
Output:
FAILURE or SVD of a rank-$r$ matrix $X\approx M$ such that $||M-X||\leq\mu$.
Initialization:
Fix an upper rank $\rho$ satisfying
$r\leq\rho\ll\sqrt{n}.$ (2.1)
Computations:
  1. 1.
    Compute an approximation of $M$ given by SVD of a matrix $M(\rho)$ of rank $\rho$.
  2. 2.
    Compute its $r$-top SVD (cf. Appendix B), which is SVD of $X:=M(\rho)_{r}$.
  3. 3.
    Estimate the norm $||E||$. If
    $||E||\leq\mu~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
    then output that SVD and $||E||$. Otherwise output FAILURE.
Next we prove bound (1.3).
Theorem 2.3.
The error norm $||E||=||M-X||$ (cf. Eqn. (2.2)) satisfies
$||E||\leq\sigma_{r+1}(M)+2\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||,$ (2.3)
which turns into (1.3) for $\sigma_{r+1}(M)=||M-M_{r}||$.
Proof.
Represent $E$ of (2.2) as $M-M(\rho)+M(\rho)-M(\rho)_{r}$, apply triangle inequality, and obtain
$||E||\leq||M-M(\rho)||+||M(\rho)-M(\rho)_{r}||=\nu+\sigma_{r+1}(M(\rho)).$
Thm. 2.2 implies that $\sigma_{r+1}(M(\rho))\leq\nu+\sigma_{r+1}(M)$. Combine the latter two bounds. ∎
Remark 2.1.
Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$, but can be very pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
Remark 2.2.
[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should keep $\rho$ small enough to perform at stage 2 fast or superfast. For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
Remark 2.3.
[A simplified version.] In Sec. 5.3 we test a simplified version of Alg. 2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$. In practice, the choice of integers $\rho$ and then $r$ should adaptive, towards minimizing the error norms of the resulting rank-$\rho$ and rank-$r$ approximations of $M$.
3 LRA by means of random sampling with sparse multipliers
The random sampling LRA algorithms of [11, 33] are close to optimal with a high probability but are not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT. In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations to superfast level by using sparse test matrices, e.g., Ultra-Sparse Rademacher test matrices [33, Sec. 3.9], [3], which depend on a positive integer parameter $s$, such that a test matrices are sparse and support superfast LRA for smaller values $s$. It is a challenge to support our VLRA with these algorithms and test matrices, but in this paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix D and also supporting superfast VLRA. For other candidate sparse test matrices see [11, Remark 4.6], [7, Sec. 5.1.8].
According to [13], such acceleration tends to make the output LRAs somewhat less reliable, and indeed in our tests LRAs computed at stage 1 of Alg. 2.1 tended to be cruder with abridged SRHT test matrices than with Gaussian random ones. Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec. 1.3. We allow at most $k$ refinement steps for a fixed tolerance $k$. As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
Algorithm 4.1.
Iterative refinement of LRA of an $(r,\rho)$-matrix.
Input and OUTPUT
as in Alg. 2.1, except that the input of Alg. 4.1 additionally includes a positive integer $k$, which bounds the number of recursive steps allowed, and an Alg. $\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
Initialization:
Write $i:=0$, $X_{0}:=0$, and $E_{0}:=M$ and fix an upper rank $\rho$ satisfying
$2r\leq\rho\ll n.$ (4.1)
Computations:
  1. 1.
    Compute an approximation of $E_{i}$ by SVD of a matrix $E_{i}(\rho)$ of rank $\rho$. [This is step 1 of Alg. 2.1 applied to $E_{i}$ rather than $M$.]
  2. 2.
    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$ by applying Alg. $\mathbb{S}$, performed with a higher precision. [This updates approximation $X_{i}$ to $M$.]
  3. 3.
    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This compresses $M_{i}(\rho)$.]
  4. 4.
    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if it satisfies the tolerance bound, that is, if the $\mu_{i}:=\mu_{i}(M,\rho)\leq\mu$.
    Otherwise output FAILURE if $i=k$.
  5. 5.
    Otherwise, update the error matrix $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1 (to resume the iteration).
Remark 4.1.
[The computational cost bound, the choice of an upper rank, and variations of Alg. 4.1.] As in Alg. 2.1 maximize the upper rank $\rho$ to decrease the output error norm, being only limited by the assumption that the overall computational cost is dominated at stage 1 (cf. (4.1)). We can allow an upper rank $\rho$ to vary as $i$ varies as long that it stays much lower than $n$ for all $i$. In Sec. 5.3 we test a variant of Alg. 4.1 where we fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
In the following extension of Thm. 2.3 we estimate the output error norm of Alg. 4.1.
Theorem 4.1.
For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,k-1.$ Then it holds that
$\mu_{i+1}:=||E_{i+1}||\leq\sigma_{r+1}(M)+2\nu_{i}.$ (4.2)
Proof.
Recall from Alg. 4.1 that
$E_{i+1}:=M-X_{i+1}~{}{\rm for}~{}X_{i+1}:=(X_{i}+E_{i}(\rho))_{r}.$
Hence
$||E_{i+1}||\leq||M-M_{i}(\rho)||+||M_{i}(\rho)-X_{i+1}||=\nu_{i}+\sigma_{r+1}(M_{i}(\rho)).$
Thm. 2.2 implies that
$\sigma_{r+1}(M_{i}(\rho))\leq\sigma_{r+1}(M)+\nu_{i}.$
Combine the latter two bounds. ∎
Remark 4.2.
The upper bound (4.2) on the error norm is within $2\nu_{i}$ from the lower bound $\sigma_{r+1}(M)$, attained where $M_{i}(\rho)=M_{\rho}$.
Next we specify how compression limits the accuracy of the output LRAs.
Lemma 4.1.
For any $m\times n$ matrix $X$ and $m\geq n$ write $\sigma_{j}(X):=\infty$ for $j<1$ and $\sigma_{j}(X):=0$ for $j>n$. Then
[…]
Proof.
Recall that
$E_{i}=M-X_{i}~{}{\rm where}~{}\operatorname{rank}(X_{i})\leq r~{}{\rm for~{}all}~{}i.$
Furthermore,
$\sigma_{j+1}(E_{i})=\min_{Y:~{}\operatorname{rank}(Y)\leq j}||E_{i}-Y||$
by virtue of Thm. 2.1.
Substitute $E_{i}-Y=M-(Y+X_{i})$ where $\operatorname{rank}(Y+X_{i})\leq j+r$ if $\operatorname{rank}(Y)\leq j$ and obtain
$\sigma_{j+1}(E_{i})\geq\min_{Z:~{}\operatorname{rank}(Z)\leq j+r}||M-Z||=\sigma_{j+r+1}(M).$
Similarly deduce that $\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$. ∎
Corollary 4.1.
The singular values $\sigma_{\rho+1}(E_{i})$ satisfy
$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,k-1.$ (4.3)
Remark 4.3.
[Error norm bounds.] In good accordance with the corollary, in our tests with $(r,\rho)$-matrices the output accuracy of Alg. 2.1 was higher than that of Alg. 4.1 provided that the upper rank $\rho$ in Alg. 2.1 was twice as large as in Alg. 4.1. The test results were mixed for input matrices with flat spectra.
5 Numerical experiments
In this section we cover the tests of our LRA algorithms on synthetic and real-world matrices. We implemented the algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory and called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
5.1 Input Matrices
[…]
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [11, Sec. 7.1] for more details).
Fig. 1 shows the distribution of the top 50 singular values of these three matrices.
We set the target ranks $r$ equal to 20 for Shaw, 45 for Gravity, and 11 for SLP matrices.
Figure 1: Spectra of singular values of some real-world matrices
We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [34, Sec. 7.3.2].
The $10738\times 5001$ matrix StreamVel is involved in numerical simulation (DNS) on a coarse mesh of the 2D Navier–Stokes equations for a low-Reynolds number flow around a cylinder.
The $19264\times 7305$ matrix MinTemp tabulates meteorological variables at weather stations across the northeastern United States on days during the years 1981–2016.
The $2000\times 2000$ matrix MaxCut provides a high-accuracy solution to the MaxCut SDP for a sparse graph [9].
We let the target rank $r$ be 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2.
Synthetic Input Matrices.
We generated random synthetic $1024\times 1024$ input matrices of five classes, two of them – with fast and slowly decaying spectra – as the products $U\Sigma V^{T}$, with the matrices $U$ and $V$ of the left and right singular vectors of a Gaussian random matrix, respectively.
By letting $\Sigma=\textrm{diag}(v)$, $v_{i}=1$ for $i=1,2,3,\dots,20$, $v_{i}=\frac{1}{2^{i-20}}$ for $i=21,\dots,100$, and $v_{i}=0$ for $i>100$, we arrived at the matrices with fast decaying spectra.
By letting $\Sigma=\textrm{diag}(u)$, $u_{i}=1$ for $i=1,2,\dots,20$, and $u_{i}=\frac{1}{(1+i-20)^{2}}$ for $i>20$, we arrived at the matrices with slowly decaying spectra.
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [34, Sec. 7.3.1], which involved effective rank parameter $R=20$.
(i) Low-rank + noise matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$ for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, depend on a parameter $\xi$ and become Low Rank Low Noise matrix for $\xi=10^{-4}$, the Low Rank Med Noise matrix for $\xi=10^{-2}$, and the Low Rank High Noise matrix for $\xi=10^{-1}$.
(ii) Polynomial Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p})$ are Poly Decay
Slow
for $p=0.5$, Poly Decay Med for $p=1$, and Poly Decay Fast for $p=2$.
(iii) Exponential Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q})$ are Exp Decay
Slow
for $q=0.01$, Exp Decay Med for $q=0.1$, and Exp Decay Fast for $q=0.5$.
5.2 Two options at stages 1 of Algs. 2.1 and 4.1
In our tests of Algs. 2.1 and 4.1 we used the following two options at their stage 1 for computations of a crude rank-$\rho$ approximation of a matrix $M$.
(i) Fast (although not superfast) randomized algorithm of [33] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{C}^{(2\rho)\times 1024}$ and $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
5.3 Test results for Algs. 2.1 and 4.1
For Alg. 2.1 we output the relative error norms $\text{err}:=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$ for input matrices $M$, a target rank $r$, the $r$-truncation $M_{r}$ and $M(\rho)_{r}$, and the upper rank $\rho$, set to $2r,3r,4r$, $5r$ (cf. Remark 2.3).
[…]
Our tables display the mean of the relative error norms observed in these tests.
For Alg. 2.1 we also display the standard deviation.
For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg. 2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg. 4.1 where TYPE stands for an input matrix type. We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT test matrices, which were more or less in the same range for both matrix types. All our tests with the synthetic input matrices from [34, Sec. 7.3.1] wit abridged SRHT test matrices have failed, that is, have output large error norms. In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random test matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
Alg. 2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT test matrices. This could have been predicted because the spectrum of the singular values $\sigma_{j}(M)$ of these matrices $M$ is flat for $j>20$ (decreasing roughly from $10^{-13}$ to between $10^{-15}$ and $10^{-16}$); in Table 5.1 we display the test results for the matrix Shaw and target rank $r=19$, which are about as favorable as for Gravity and Slow Decay.
In both cases of success and failure of both algorithms, our test results were in good accordance with our formal study. The relative error norms output by both Algs. 2.1 and 4.1 applied to $(r,\rho)$-matrices have consistently decreased and frequently to the optimal value 1 (up to the measurement errors) as we increased the upper rank $\rho$ for Alg. 2.1 and in the transition from the first to the second iteration of Alg. 4.1. In all tests except for High Noise, Slow Decay, and MinTemp matrices already err2 decreased below 1.1. The third iteration decreased the error norm much less or even slightly increased it.
For $(r,\rho)$-matrices the error norm bounds of Alg. 2.1 for $\rho=4r$ were consistently smaller than those of Alg. 4.1 in all three iterations. This was not the case just for the matrices with flat spectra such as the matrices Shaw for $r=20$, High Noise, and Exp Decay Slow.
Input Matrix Multiplier $\rho$
$2r$ $3r$ $4r$ $5r$
Gravity Abr. SRHT 1.000 $\pm$ 1.036E-05 1.000 $\pm$ 7.815E-06 1.000 $\pm$ 7.888E-06 1.000 $\pm$ 6.683E-06
Gaussian 1.000 $\pm$ 7.495E-06 1.000 $\pm$ 5.211E-06 1.000 $\pm$ 7.214E-06 1.000 $\pm$ 5.432E-06
SLP Abr. SRHT 1.970 $\pm$ 4.083 1.000 $\pm$ 0.002 1.000 $\pm$ 1.018E-07 1.000 $\pm$ 1.362E-10
Gaussian 1.001 $\pm$ 0.0008160 1.000 $\pm$ 1.528E-07 1.000 $\pm$ 1.075E-10 1.000 $\pm$ 6.350E-14
Shaw, $r=19$ Abr. SRHT 1.000 $\pm$ 1.778E-06 1.000 $\pm$ 1.359E-06 1.000 $\pm$ 1.665E-06 1.000 $\pm$ 1.441E-06
Gaussian 1.000 $\pm$ 2.256E-06 1.000 $\pm$ 1.685E-06 1.000 $\pm$ 1.260E-06 1.000 $\pm$ 1.427E-06
Fast Decay Abr. SRHT 1.000 $\pm$ 3.086E-12 1.000 $\pm$ 3.590E-16 1.000 $\pm$ 2.470E-16 1.000 $\pm$ 2.617E-16
Gaussian 1.000 $\pm$ 6.651E-12 1.000 $\pm$ 5.117E-16 1.000 $\pm$ 4.552E-16 1.000 $\pm$ 4.590E-16
Slow Decay Abr. SRHT 1.000 $\pm$ 3.604E-05 1.000 $\pm$ 1.244E-06 1.000 $\pm$ 1.543E-07 1.000 $\pm$ 3.754E-08
Gaussian 1.000 $\pm$ 4.130E-05 1.000 $\pm$ 1.380E-06 1.000 $\pm$ 1.783E-07 1.000 $\pm$ 3.708E-08
Table 5.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, and synthetic matrices Fast Decay and Slow Decay, with fast and slowly decaying spectra.
Input Matrix Multiplier 1st Itr. 2nd Itr. 3rd Itr.
Fast Decay Abr. SRHT 3.155000 1.0000 1.0000
Gaussian 3.1202 1.0000 1.0000
Slow Decay Abr. SRHT 5.0468 1.0003 1.0001
Gaussian 5.0755 1.0002 1.0001
Shaw for $r=20$ Abr. SRHT 2.8820E+01 1.0983 1.1225
Gaussian 1.8235E+01 1.1517 1.1189
Gravity Abr. SRHT 1.5762E+01 1.0000 1.0000
Gaussian 1.2917E+01 1.0000 1.0000
SLP Abr. SRHT 1.0931E+02 1.0014 1.0000
Gaussian 5.2205 1.0000 1.0000
Table 5.2: Test results for Alg. 4.1 on Fast Decay, Slow Decay, Shaw, Gravity, and SLP matrices.
Input Matrix $\rho$
$2r$ $3r$ $4r$ $5r$
Low Rank Low Noise 1.0416 $\pm$ 8.9977E-02 1.$0$$\pm$ 2.1197E-06 1.$0$$\pm$ 2.4039E-06 1.$0$$\pm$ 2.2834E-06
Low Rank Med Noise 1.4335 $\pm$ 1.7048E-01 1.0382 $\pm$ 3.1809E-02 1.0057 $\pm$ 1.4442E-03 1.0026 $\pm$ 5.8650E-04
Low Rank High Noise 5.6972 $\pm$ 8.6182E-01 4.8401 $\pm$ 4.3819E-01 4.0328 $\pm$ 2.3493E-01 3.7893 $\pm$ 2.1626E-01
Poly Decay Slow 2.0588 $\pm$ 1.8783E-01 1.6525 $\pm$ 2.0027E-01 1.3617 $\pm$ 8.6188E-02 1.2062 $\pm$ 8.6352E-02
Poly Decay Med 1.5384 $\pm$ 2.1907E-01 1.0315 $\pm$ 2.6585E-02 1.0028 $\pm$ 1.0190E-03 1.0009 $\pm$ 4.5652E-04
Poly Decay Fast 1.3133 $\pm$ 1.5431E-01 1.0001 $\pm$ 1.3887E-04 1.$0$$\pm$ 3.7489E-06 1.$0$$\pm$ 2.4495E-07
Exp Decay Slow 2.8587 $\pm$ 3.3389E-01 2.2772 $\pm$ 2.0481E-01 1.8244 $\pm$ 1.0970E-01 1.5721 $\pm$ 1.0528E-01
Exp Decay Med 1.5576 $\pm$ 1.2324E-01 1.0414 $\pm$ 4.9018E-02 1.0001 $\pm$ 9.0188E-05 1.$0$$\pm$ 3.7953E-07
Exp Decay Fast 1.3121 $\pm$ 1.4989E-01 1.$0$$\pm$ 6.4663E-11 1.$0$$\pm$ 3.6020E-16 1.$0$$\pm$ 3.0986E-16
Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian test matrices.
Input Matrix 1st Itr. 2nd Itr. 3rd Itr.
Low Rank Low Noise 1.3940 1.0000 1.0000
Low Rank Med Noise 1.4752 1.0386 1.0375
Low Rank High Noise 1.4507 1.5135 1.5332
Poly Decay Slow 1.5920 1.4154 1.4073
Poly Decay Med 1.5569 1.0345 1.0306
Poly Decay Fast 1.3784 1.0001 1.0002
Exp Decay Slow 1.5202 1.4956 1.4750
Exp Decay Med 1.4946 1.0115 1.0164
Exp Decay Fast 1.4235 1.0000 1.0000
Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian multipliers.
Input Multiplier $\rho$
Matrix Type $2r$ $3r$ $4r$ $5r$
StreamVel Abr. SRHT 1.5840 $\pm$ 1.2583 1.4736 $\pm$ 1.4538 1.0022 $\pm$ 2.6478E-03 1.0002 $\pm$ 2.7733E-04
Gaussian 1.0932 $\pm$ 5.8887E-02 1.0064 $\pm$ 2.9809E-03 1.0013 $\pm$ 1.4141E-03 1.0003 $\pm$ 3.1762E-04
MinTemp Abr. SRHT 2.4575 $\pm$ 2.5748E-01 1.6980 $\pm$ 2.3998E-01 1.2919 $\pm$ 1.3606E-01 1.1181 $\pm$ 9.7415E-02
Gaussian 2.4393 $\pm$ 3.0823E-01 1.7216 $\pm$ 2.2510E-01 1.3793 $\pm$ 7.6629E-02 1.1238 $\pm$ 8.3331E-02
MaxCut Abr. SRHT 1.0024 $\pm$ 1.3630E-03 1.0005 $\pm$ 1.6875E-04 1.0002 $\pm$ 6.8092E-05 1.0001 $\pm$ 2.9040E-05
Gaussian 1.0026 $\pm$ 1.3856E-03 1.0005 $\pm$ 2.1129E-04 1.0002 $\pm$ 6.6260E-05 1.0001 $\pm$ 2.4207E-05
Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [34, Sec. 7.3.2]
Input Matrix Multiplier Type 1st Itr. 2nd Itr. 3rd Itr.
StreamVel Abr. SRHT 2.9700 1.0739 1.0300
Gaussian 2.8724 1.0397 1.0177
MinTemp Abr. SRHT 2.3340 1.5961 1.5079
Gaussian 2.3120 1.6410 1.4950
MaxCut Abr. SRHT 3.6200E+01 1.0312 1.0182
Gaussian 2.6202E+01 1.0302 1.0176
Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [34, Sec. 7.3.2].
6 Conclusions
Rank-$r$ approximation of an $m\times n$ matrix $M$ is only interesting if there is a gap in the spectrum of its top (largest) $r+1$ singular values $\sigma_{j}(M)$. For the $(r,\rho)$-matrices $M$ having also a gap in the next segment of their spectra – from $(r+1)$st to $(\rho+1)$st largest singular values – randomized oversampling algorithms output quite accurate LRA.
We extended these observations towards devising superfast (sublinear cost) variation of LRA algorithms based on random sampling and oversampling with sparse test matrices and to improving accuracy of the resulting LRA in the case of $(r,\rho)$-matrices $M$ where $r<\rho\ll\sqrt{n}$. We provided some formal and experimental support for the efficiency of the resulting LRA algorithms.
There are various natural extensions of our study, which can be practically competitive. E.g., one can devise and test similar variations of the TYUC and GN algorithms with sparse test matrices such as Ultra-Sparse Rademacher test matrices [33, Sec. 3.9] or sparsified random Givens rotations by Rokhlin and Tygert (see [11, Remark 4.6]).666[26, 27, 28] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes.
Our study in Secs. 4 and 5 has showed no benefits for using iterative refinement of LRA versus oversampling so far, but one can keep comparing their efficiency.
Appendix
Appendix A Small families of matrices that are hard for superfast LRA
Any superfast algorithm fails to compute a close LRA of the following small families of matrices.
Example A.1.
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Likewise, we can verify that any superfast randomized LRA algorithms fail fails on those matrix families with error probability not close to 0.
Remark A.1.
How representative are the above matrix families of hard inputs? The matrices of these families represent data singularities and are not relevant to a large class of input matrices representing regular processes or, say, smooth surfaces. Furthermore, Example A.1 does not apply to various important special classes of input matrices such as symmetric or diagonally dominant.
Appendix B Computation of $r$-top SVD of a matrix product
For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [22, 30]) based on computation of the SVDs of $A$ and $B$.
Algorithm B.1.
[Computation of an $r$-top SVD of a matrix product.]
Input:
Four integers $r$, $k$, $m$, and $n$ such that $0<r\leq k\leq\min\{m,n\}$ and two matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$.
Output:
Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{C}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
Computations:
  1. 1.
    Compute SVDs $A=U_{A}\Sigma_{A}V^{*}_{A}$ and $B=U_{B}\Sigma_{B}V^{*}_{B}$ where $U_{B}\in\mathbb{C}^{m\times k}$, $V_{B}^{*}\in\mathbb{C}^{k\times n}$, and $\Sigma_{A},V^{*}_{A},U_{B},\Sigma_{B}\in\mathbb{C}^{k\times k}$.
  2. 2.
    Compute $k\times k$ matrices $W=\Sigma_{A}V^{*}_{A}U_{B}\Sigma_{B}$, $U_{W}$, $\Sigma_{W}$, and $V^{*}_{W}$ such that $W=U_{W}\Sigma_{W}V_{W}^{*}$ is SVD, $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$, and $\Sigma=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ and $\Sigma^{\prime}=\operatorname{diag}(\sigma_{j})_{j=r+1}^{k}$ are the matrices of the $r$ top (largest) and the $k-r$ trailing (smallest) singular values of the matrix $W$, respectively. Output the matrix $\Sigma$.
  3. 3.
    Compute and output the matrices $U$ and $V$ made up of the first $r$ columns of the matrices $U_{A}U_{W}$ and $V_{B}V_{W}$, respectively.
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [7, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
Appendix C CUR decomposition, CUR LRA, and its link to top SVD
C.1 CUR decomposition and CUR approximation
Next we briefly revisit CUR LRA, which is particularly memory efficient. For five integers $k$, $\ell$, $m$, $n$, and $r$ satisfying (1.2), an $m\times n$ matrix $M$, and two sets $\mathcal{I}\subseteq\{1,\dots,m\}$ and $\mathcal{J}\subseteq\{1,\dots,n\}$ of cardinalities $k$ and $\ell$, respectively, define the submatrices
$R:=M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},C:=M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{}{\rm and}~{}G:=M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}$
of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$, although we recall notation $G$ for nucleus and CGR for the factorization in [21], originated with the decomposition itself in the 1990s [8]. In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
Given a $k\times\ell$ generator submatrix $G$ of $M$, we can compute CUR approximation of $M$ by using $k\ell$ scalars and $O(k\ell\min\{k,\ell\})$ flops.
Theorem C.1.
(See [28].) For matrices $C$, $U$, and $R$ defined above, $M=CUR$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large.
C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
The algorithm of Appendix B transforms a CUR decomposition into a top SVD. The next algorithm makes converse transition. It uses $(m+n+r)r$ scalars and order of $(m+n)r^{2}$ flops to compute CUR decomposition having an $r\times r$ generator for a matrix given with its $r$-top SVD, for $r\leq n\leq m$. By combining this algorithm with that of Appendix B we can transform an LRA of a matrix into its CUR LRA at a low computational cost.
Algorithm C.1.
[From $r$-top SVD to CNR decomposition.] [^7] We denote nucleus by $N$ rather than $U$ or $G$ to avoid conflicts with the factor $U$ in SVD and generator $G$ of CUR of Sec. C.1.
Input:
Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{C}^{m\times n}$, $\Sigma\in\mathbb{C}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{C}^{m\times r}$, and $V\in\mathbb{C}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
Output:
Three matrices $C\in\mathbb{C}^{m\times r}$, $N\in\mathbb{C}^{r\times r}$, and $R\in\mathbb{C}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
$M=CNR.$
Computations:
By applying to the matrices $U$ and $V$ the algorithms of [6] or [23] compute two nonsingular $r\times r$ submatrices $U_{\mathcal{I},:}$ and $V^{*}_{:,\mathcal{J}}$.
Compute and output the CUR factors $C=U\Sigma V^{*}_{:,\mathcal{J}}$, $R=U_{\mathcal{I},:}\Sigma V^{*}$, and $N:=(V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1}$.
Correctness verification. Substitute the expressions for $C$, $N$ and $R$ and obtain $CNR=(U\Sigma V^{*}_{:,\mathcal{J}})((V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1})(U_{\mathcal{I},:}\Sigma V^{*})=U\Sigma V^{*}$.
Cost bounds. The algorithm uses $O(mr^{2})$ flops and $(m+n+r)r$ scalars; these cost bounds are sublinear where $r^{2}\ll n$ and are dominated at stage 2.
Let us also estimate the norm of the nucleus $||N||$. Observe that $\operatorname{rank}(V_{:,\mathcal{J}})=\operatorname{rank}(\Sigma)=\operatorname{rank}(U_{\mathcal{I},:})=r$ and $||N||\leq||V_{:,\mathcal{J}}^{-1}||~{}||\Sigma^{-1}||~{}||U_{\mathcal{I},:}^{-1}||$ and recall that $||\Sigma^{-1}||=||M^{+}||=1/\sigma_{r}(M)$.
Write $t_{q,s,h}^{2}:=(q-s)sh^{2}+1$, allow any choice of $h>1$, say, $h=1.1$, and then recall that $||U_{\mathcal{I},:}^{+}||\leq t_{m,k,h}^{a}$, $||(V_{:,\mathcal{J}}^{*})^{+}||\leq t_{n,\ell,h}^{a}$, and consequently
$||N||\leq t_{m,r,h}^{a}t_{n,r,h}^{a}/\sigma_{r}(M)$
where $a=1$ if we apply the algorithms of [6] at stage and $a=2$ if we apply those of [23].
Appendix D Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in our tests for Algs. 2.1 and 4.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [17, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (D.1)
For any fixed pair of $d$ and $i$, each of the matrices $H_{d,i}$ is orthogonal up to scaling and has $2^{d}$ nonzero entries in every row and column; we can compute the product $MH$ for an $n\times k$ submatrix $H$ of $H_{d,d}$ by using less than $km2^{d}$ additions and subtractions.
Now define the $d$-Abridged Scaled and Permuted Hadamard matrices, $PDH_{d,d}$, where $P$ is a random sampling matrix and $D$ is the matrix of random integer diagonal scaling. Each random permutation or scaling contributes up to $n$ random parameters. We can involve more random parameters by applying random permutation and scaling also to some or all intermediate matrices $H_{d,i}$ for $i=0,1,\dots,d$.
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [11, Sec. 11]).
Acknowledgement: Jianlin Xia helped place our study into proper historical context.