### v17

# Iterative Refinement and Oversampling for Low Rank Approximation

Abstract
Iterative refinement is particularly popular for numerical solution of linear systems of equations. We extend it to Low Rank Approximation of a matrix (LRA) and observe close link of the resulting algorithm to oversampling techniques, commonly used in randomized LRA algorithms. We elaborate upon this link and revisit oversampling and some efficient randomized LRA algorithms. Applied with sparse sketch matrices they run significantly faster and in particular yield Very Low Rank Approximation (VLRA) at sublinear cost, using much fewer scalars and flops than the input matrix has entries. This is achieved at the price of deterioration of output accuracy, but according to our formal and empirical study subsequent oversampling improves accuracy to near-optimal level under the spectral norm for a large sub-class of matrices with fast decaying spectra of singular values.
Key Words:
low rank approximation, iterative refinement, oversampling, sparse sketches, singular values, gaps in the spectrum, sublinear cost.
2020 Math. Subject Classification:
65F55, 65N75, 65Y20, 68Q25, 68W20
1 Introduction: LRA of a matrix, its iterative refinement, link to oversampling, and sparse sketches
1.1 Brief overview
Classical iterative refinement (see pointers to the bibliography in Sec. 1.6) is particularly popular for numerical solution of linear systems of equations. We naturally extend it to Low Rank Approximation of a matrix (LRA), observe its close link to LRA with oversampling, and analyze both refinement and oversampling. Here is a brief outline of our study.
Let an LRA algorithm $\mathbb{A}=\mathbb{A}_{r}$ be applied to an $m\times n$ matrix $E_{0}:=M$ and let it output a rank-$r$ approximation $Y_{0}:=X_{0}:=\mathbb{A}(E_{0})$ of $M$.  [^1] Here and hereafter “rank-$r$ approximation” means “rank at most $r$ approximation”. Apply the algorithm $\mathbb{A}$ to the error matrix $E_{1}=E_{0}-X_{0}$ and add $X_{1}:=\mathbb{A}(E_{1})$ to $X_{0}$ to obtain improved approximation $Y_{1}:=X_{0}+X_{1}$ to $M$. By repeating this step recursively $h$ times, compute the matrices $X_{i}:=\mathbb{A}(E_{i})$, $Y_{i}:=\mathbb{A}(Y_{i-1}+X_{i})$, and $E_{i+1}:=E_{i}-X_{i}$, for $i=1,2,\dots,h$. Finally output the matrix $\mathbb{A}(Y_{h})$ as a rank-$r$ approximation of $M$. The compression of higher rank approximation of a matrix into its lower rank approximation also occurs in the customary random oversampling for LRA.
We revisit random sampling LRA algorithms with oversampling, accelerate them by using sparse sketch matrices and in particular yield Very Low Rank Approximation (VLRA) at sublinear cost. We achieve this at the price of some deterioration of the accuracy of the output LRA, but according to our formal and empirical study, subsequent oversampling improves accuracy to near-optimal under the spectral norm for a large sub-class of the matrices with fast decaying spectra of singular values. So far, our study has showed no additional benefits for LRA from application of iterative refinement versus standalone oversampling.
1.2 The State of the Art of LRA
LRA is a hot topic of Numerical Linear Algebra. We first briefly recall the State of the Art, referring to [10, 27, 15] for broader expositions, and then outline our results for LRA.
1.1.1. Representation and the power of LRA. Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors. [^2] The sketch of a matrix in [30, 27] is its product with a matrix of smaller size, that is, with a small set of vectors. In this way we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [27, Sec. 1.3]). Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [29], that is, admit LRA, with which we can operate efficiently, even where $M$ has enormous size.
Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r<n$ or a fixed target range for such a rank, although actually a proper choice of such targets is a challenge (see Remark 1.1). Then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
$X\in\mathbb{C}^{m\times k},Y\in\mathbb{C}^{k\times\ell},~{}Z\in\mathbb{C}^{\ell\times n},$ (1.1)
$k\leq m,~{}\ell\leq n,~{}\min\{k,\ell\}=r,$ (1.2)
and the spectral and/or Frobenius norm of the matrix $M-XYZ$ is small in context, say, close to optimal.
1.1.2. LRA from top SVD; SVD of an LRA. By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, turn SVD of $M$ into its $r$-top SVD and turn $M$ into its $r$-truncation $M_{r}$, of rank $r$. By virtue of the Eckart-Young-Mirsky theorem [6, Thm. 2.4.8], $r$-top SVD defines an optimal rank-$r$ approximation of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [6, Fig. 8.6.1]). This is too expensive for LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ by using just $(m+k)k+(n+\ell)\ell$ scalars and $O(mk^{2}+n\ell^{2})$ flops (see Appendix B).
1.1.3. LRA via random sampling and oversampling. Random sampling with oversampling enables fast computation of close LRA (see [2, 14, 10, 30, 27, 16], and the references therein). For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [10] successively generates a random sketch matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
Remark 1.1.
“In practice, the target rank $k$ is rarely known in advance. Randomized algorithms are usually implemented in an adaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [10, Sec. 4.2]).
In the best studied case the sketch matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables. In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [10, Thm. 10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops. The integer $p$ and the size of the matrix $H$ explode as $\epsilon\mapsto 0$, but LRA for $\epsilon$ as large as 1/2 or 1 is still valuable.
In other LRA algorithms (e.g., in [4, 2, 27, 16]) $M$ is both pre- and post-multiplied by random sketch matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$, respectively. In particular, the TYUC algorithm of [27], modifying those of [2, Thm. 4.7] and [2, Thm. 4.3, display 1], and the generalized Nyström approximation GN of [16] first compute the products $FM$ and $MH$ for random sketch matrices $F$ and $H$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$. In the case of random Gaussian sketch matrices $F$ and $H$ the expected Frobenius error norm of the output LRA is within a factor of 2 from optimal (namely, it does not exceed the optimal one for rank-$\frac{k}{2}$ approximation). For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [16]) and $O(\ell mn)$ flops.
The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [10, Secs. 10 and 11], [30, Ch. 6.2], [5, Thm. 3.4]).
1.3 Acceleration of randomized LRA by using sparse sketches
Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$. We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$.
We rely on combining three simple observations, namely, that (i) random sampling with oversampling is the bottleneck stage of HMT, TYUC, and GN algorithms, (ii) which we can greatly accelerate by using sparse sketch matrices (in particular abridged SRHT matrices of Appendix C), and (iii) while the accuracy of the LRA output in step (ii) tends to deteriorate with sparse sketch matrices, we readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$.
Our LRA Alg. 2.1 of Sec. 2 relies on these observations. It first computes (1) a crude rank-$\rho$ approximation $M(\rho)$ of $M$; then (2) $r$-top SVD of $M(\rho)$, output as rank-$r$ approximation of $M$.
LRA via random sampling with oversampling fits this outline.
Apart from the stage of multiplication by sketch matrices, performed at a dominated cost, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$. Unlike the number of scalars, the bound on the number of flops decreases greatly where $k+p\ll n$ and $k\leq\ell\ll n$ for HMT and for TYUC and GN algorithms, respectively. Here and hereafter $a\ll b$ means that $b$ exceeds $a$ greatly in context, e.g., $a=o(b)$ as $b\rightarrow\infty$. In particular the algorithms HMT and TYUC/GN run at sublinear cost, that is, use much fewer than $mn$ scalars and flops, where we seek Very Low Rank Approximation (VLRA), namely, where $k+p\ll\sqrt{n}$ or $kl\ll\sqrt{n}$, respectively.
Alg. B.1 of Appendix B performs Stage (2) by using about as many scalars and $O(m\ell^{2})$ flops. We can decrease this bound to $O(mk\ell)$ (which is sublinear for $kl\ll\sqrt{n}$) by applying TUYC or GN algorithms to compute near-optimal rank-$r$ approximation of $M(\rho)$ rather than its $r$-top SVD at stage (2) of Alg. 2.1.
1.4 Accuracy of the output LRA
By virtue of Thm. 2.3, the spectral norm $||\cdot||:=||\cdot||_{2}$ of the output error matrix of Alg. 2.1 satisfies
$||M-M(\rho)_{r}||\leq||M-M_{r}||+2||M-M(\rho)||,$ (1.3)
that is, Alg. 2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$. For $(r,\rho)$-matrices, this factor can be large only if the computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails. It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse sketch matrices tend to be unsatisfactory as LRA [11], [27, Sec. 3.9], but in our numerical tests in Sec. 5 accuracy of the output matrices of stage (1) of Alg. 2.1 tended to be raised at stage (2) to near-optimal level for a large sub-class of $(r,\rho)$-matrices $M$.
1.5 Iterative LRA
In Sec. 4 we apply our LRA algorithm (or we can apply any efficient LRA algorithm instead) to the output error matrix of a crude LRA computed at stage (1), to decrease its norm. Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs. We reduce the algorithm essentially to recursive applications of Alg. 2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance. We counter this deficiency by recursively compressing the sums into rank-$r$ matrices, although this limits output accuracy of computed LRAs. The resulting iterative LRA algorithm is a natural extension to LRA of an efficient popular technique of matrix computations; we specify, analyze, and test it in Secs. 4 and 5.
Our upper estimates for the accuracy of both Alg. 2.1 and iterative refinement are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
In our tests, the output accuracy of our iterative LRA applied to $(r,\rho)$-matrices improved significantly at the second iteration, little improved afterwards, and overall error norm was never less than with Alg. 2.1 for the upper rank $\rho=4r$.
1.6 Related works
Our iterative LRA extends the algorithms of [13, 19], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [26, Secs. 3.3.4 and 4.2.5], [9, Ch. 12 and Sec. 20.5], [6, Secs. 3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [26, page 223 and the references on page 225]). Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [18, 24]. Such compression is commonly used for LRA with oversampling. For sparse test matrices we used abridged SRHT matrices studied in the papers [20, 21, 22], which in turn extends the earlier study of Gaussian elimination without pivoting in [23, 25].
Unlike our presentation in Secs. 1.3 and 1.5 as well as in Secs. 2 and 4, we actually first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.
1.7 Organization of our paper
We present our accelerated oversampling LRA algorithm in Sec. 2, discuss fast computation of a crude LRA in Sec. 3, present our iterative LRA algorithm in Sec. 4, and cover our numerical experiments in Sec. 5. We devote our short Sec. 6 to conclusions. In Appendix A we describe some small families of matrices for which any LRA algorithm fails unless it access all entries of an input matrix. In Appendix B we fast compute $r$-top SVD of an LRA of a low-rank matrix. In Appendix C we generate abridged SRHT matrices, whose incorporation as sketch matrices makes random sampling LRA run faster.
2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
We begin with some background materials.
Hereafter $||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Remark 2.2. In particular we only use Eckart-Young’s (rather than Mirski’s) theorem:
Theorem 2.1.
$\sigma_{r+1}(M)=||M_{r}-M||=\min_{X:~{}\operatorname{rank}(X)\leq r}||M-X||$.
Theorem 2.2.
[6, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
$|\sigma_{j}(M+E)-\sigma_{j}(M)|\leq||E||~{}{\rm for}~{}j=1,\dots,\min\{m,n\}.$
In the following algorithm we elaborate upon the outline of Sec. 1.3 assuming that we are given black box subroutines that output a crude LRA and an estimate for a spectral norm of a matrix.
Algorithm 2.1.
LRA of $(r,\rho)$-matrices.
Input:
An $m\times n$ matrix $M$, for $n\leq m$, given implicitly by an oracle (black box subroutine) for its multiplication by a vector (cf. Sec. 1.1.1), a target rank $r\ll n$, and a tolerance $\mu$ to the output error norm.
Output:
FAILURE or SVD of a rank-$r$ matrix $X\approx M$ such that $||M-X||\leq\mu$.
Initialization:
Fix an upper rank $\rho$ satisfying
$r\leq\rho\ll\sqrt{n}.$ (2.1)
Computations:
  1. 1.
    Compute an approximation of $M$ given by SVD of a matrix $M(\rho)$ of rank $\rho$.
  2. 2.
    Compute its $r$-top SVD (cf. [19, Appendix B]), which is SVD of $X:=M(\rho)_{r}$.
  3. 3.
    Estimate the norm $||E||$. If
    $||E||\leq\mu~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
    then output that SVD and $||E||$. Otherwise output FAILURE.
Next we prove bound (1.3).
Theorem 2.3.
The error norm $||E||=||M-X||$ (cf. Eqn. (2.2)) satisfies
$||E||\leq\sigma_{r+1}(M)+2\nu~{}{\rm where}~{}\nu=\nu(M,\rho):=||M-M(\rho)||,$ (2.3)
which turns into (1.3) for $\sigma_{r+1}(M)=||M-M_{r}||$.
Proof.
Represent $E$ of (2.2) as $M-M(\rho)+M(\rho)-M(\rho)_{r}$, apply triangle inequality, and obtain
$||E||\leq||M-M(\rho)||+||M(\rho)-M(\rho)_{r}||=\nu+\sigma_{r+1}(M(\rho)).$
Thm. 2.2 implies that $\sigma_{r+1}(M(\rho))\leq\nu+\sigma_{r+1}(M)$. Combine the latter two bounds. ∎
Remark 2.1.
Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$ but can be quite pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
Remark 2.2.
[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should choose it smaller to perform stage 2 faster. For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
Remark 2.3.
[A simplified version.] In Sec. 5.3 we test a simplified version of Alg. 2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$. In practice, the choice of integers $\rho$ and then $r$ should be adapted towards minimizing the error norms of the computed rank-$\rho$ and rank-$r$ approximations of $M$.
3 LRA by means of random sampling with sparse sketch matrices
The random sampling LRA algorithms of [10, 27] are close to optimal with a high probability under various customary choices of dense sketch matrices such as Gaussian, Rademacher, SRHT, and SRFT. In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations by using sparse sketch matrices, such as Ultra-Sparse Rademacher sketch matrices [27, Sec. 3.9], [2], which depend on a positive integer parameter $s$, such that sketch matrices are sparse and support faster LRA for smaller values $s$. In the present paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix C and also supporting faster VLRA. For other candidate sparse sketch matrices see [10, Remark 4.6], [6, Sec. 5.1.8].
According to [11], such acceleration tends to make the output LRAs somewhat less reliable, and in our tests LRAs computed at stage 1 of Alg. 2.1 tended to be cruder with abridged SRHT sketch matrices than with Gaussian random ones. Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec. 1.5. As a basis we apply Alg. 2.1, but other efficient LRA algorithms can be applied instead of it. We allow at most $h$ refinement steps for a fixed tolerance $h$. As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
Algorithm 4.1.
Iterative refinement of LRA of an $(r,\rho)$-matrix.
Input and OUTPUT
as in Alg. 2.1, except that the input of Alg. 4.1 additionally includes a positive integer $h$, which bounds the number of recursive steps allowed, and an Alg. $\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
Initialization:
Write $i:=0$, $X_{0}:=0$, and $E_{0}:=M$ and fix an upper rank $\rho$ satisfying
$2r\leq\rho\ll n.$ (4.1)
Computations:
  1. 1.
    Compute an approximation of $E_{i}$ by SVD of a matrix $E_{i}(\rho)$ of rank $\rho$. [This is step 1 of Alg. 2.1 applied to $E_{i}$ rather than $M$.]
  2. 2.
    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$ by applying Alg. $\mathbb{S}$, performed with a higher precision. [This updates approximation $X_{i}$ to $M$.]
  3. 3.
    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This enables us to compress $M_{i}(\rho)$.]
  4. 4.
    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if it satisfies the tolerance bound, that is, if the $\mu_{i}:=\mu_{i}(M,\rho)\leq\mu$.
    Otherwise output FAILURE if $i=h$.
  5. 5.
    Otherwise, update the error matrix $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1 (to resume the iteration).
Remark 4.1.
[The computational cost bound, the choice of an upper rank, and variations of Alg. 4.1.] As in Alg. 2.1 maximize the upper rank $\rho$ to decrease the output error norm, being only limited by the assumption that the overall computational cost is dominated at stage 1 (cf. (4.1)). We can allow an upper rank $\rho$ to vary as $i$ varies as long that it stays much lower than $n$ for all $i$. In Sec. 5.3 we test a variant of Alg. 4.1 where we fix $h=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
In the following extension of Thm. 2.3 we estimate the output error norm of Alg. 4.1.
Theorem 4.1.
For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,h-1.$ Then it holds that
$\mu_{i+1}:=||E_{i+1}||\leq\sigma_{r+1}(M)+2\nu_{i}.$ (4.2)
Proof.
Recall from Alg. 4.1 that
$E_{i+1}:=M-X_{i+1}~{}{\rm for}~{}X_{i+1}:=(X_{i}+E_{i}(\rho))_{r}.$
Hence
$||E_{i+1}||\leq||M-M_{i}(\rho)||+||M_{i}(\rho)-X_{i+1}||=\nu_{i}+\sigma_{r+1}(M_{i}(\rho)).$
Thm. 2.2 implies that
$\sigma_{r+1}(M_{i}(\rho))\leq\sigma_{r+1}(M)+\nu_{i}.$
Combine the latter two bounds. ∎
Remark 4.2.
The upper bound (4.2) on the error norm is within $2\nu_{i}$ from the lower bound $\sigma_{r+1}(M)$, attained where $M_{i}(\rho)=M_{\rho}$.
Next we specify how compression limits the accuracy of the output LRAs.
Lemma 4.1.
For any $m\times n$ matrix $X$ and $m\geq n$ write $\sigma_{j}(X):=\infty$ for $j<1$ and $\sigma_{j}(X):=0$ for $j>n$. Then
[…]
Proof.
Recall that
$E_{i}=M-X_{i}~{}{\rm where}~{}\operatorname{rank}(X_{i})\leq r~{}{\rm for~{}all}~{}i.$
Furthermore,
$\sigma_{j+1}(E_{i})=\min_{Y:~{}\operatorname{rank}(Y)\leq j}||E_{i}-Y||$
by virtue of Thm. 2.1.
Substitute $E_{i}-Y=M-(Y+X_{i})$ where $\operatorname{rank}(Y+X_{i})\leq j+r$ if $\operatorname{rank}(Y)\leq j$ and obtain
$\sigma_{j+1}(E_{i})\geq\min_{Z:~{}\operatorname{rank}(Z)\leq j+r}||M-Z||=\sigma_{j+r+1}(M).$
Similarly deduce that $\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$. ∎
Corollary 4.1.
The singular values $\sigma_{\rho+1}(E_{i})$ satisfy
$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,h-1.$ (4.3)
Remark 4.3.
[Error norm bounds.] In good accordance with the corollary, in our tests with $(r,\rho)$-matrices the output accuracy of Alg. 2.1 was higher than that of Alg. 4.1 provided that the upper rank $\rho$ in Alg. 2.1 was twice as large as in Alg. 4.1. The test results were mixed for input matrices with flat spectra.
5 Numerical experiments
In this section we cover the tests of our LRA algorithms on synthetic and real-world matrices. We implemented the algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory and called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
5.1 Input Matrices
[…]
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [10, Sec. 7.1] for more details).
Fig. 1 shows the distribution of the top 50 singular values of these three matrices.
We set the target ranks $r$ equal to 20 for Shaw, 45 for Gravity, and 11 for SLP matrices.
Figure 1: Spectra of singular values of some real-world matrices
We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [28, Sec. 7.3.2].
The $10738\times 5001$ matrix StreamVel is involved in numerical simulation (DNS) on a coarse mesh of the 2D Navier–Stokes equations for a low-Reynolds number flow around a cylinder.
The $19264\times 7305$ matrix MinTemp tabulates meteorological variables at weather stations across the northeastern United States on days during the years 1981–2016.
The $2000\times 2000$ matrix MaxCut provides a high-accuracy solution to the MaxCut SDP for a sparse graph [8].
We let the target rank $r$ be 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2.
Synthetic Input Matrices.
We generated random synthetic $1024\times 1024$ input matrices of five classes, two of them – with fast and slowly decaying spectra – as the products $U\Sigma V^{T}$, with the matrices $U$ and $V$ of the left and right singular vectors of a Gaussian random matrix, respectively.
By letting $\Sigma=\textrm{diag}(v)$, $v_{i}=1$ for $i=1,2,3,\dots,20$, $v_{i}=\frac{1}{2^{i-20}}$ for $i=21,\dots,100$, and $v_{i}=0$ for $i>100$, we arrived at the matrices with fast decaying spectra.
By letting $\Sigma=\textrm{diag}(u)$, $u_{i}=1$ for $i=1,2,\dots,20$, and $u_{i}=\frac{1}{(1+i-20)^{2}}$ for $i>20$, we arrived at the matrices with slowly decaying spectra.
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [28, Sec. 7.3.1], which involved effective rank parameter $R=20$.
(i) Low-rank + noise matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$ for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, depend on a parameter $\xi$ and become Low Rank Low Noise matrix for $\xi=10^{-4}$, the Low Rank Med Noise matrix for $\xi=10^{-2}$, and the Low Rank High Noise matrix for $\xi=10^{-1}$.
(ii) Polynomial Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p})$ are Poly Decay
Slow
for $p=0.5$, Poly Decay Med for $p=1$, and Poly Decay Fast for $p=2$.
(iii) Exponential Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q})$ are Exp Decay
Slow
for $q=0.01$, Exp Decay Med for $q=0.1$, and Exp Decay Fast for $q=0.5$.
5.2 Two options at stages 1 of Algs. 2.1 and 4.1
In our tests of Algs. 2.1 and 4.1 we used the following two options at their stage 1 for computations of a crude rank-$\rho$ approximation of a matrix $M$.
(i) Fast randomized algorithm of [27] with Gaussian random sketch matrices $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
(ii) Its accelerated randomized variant where abridged (length 3) SRHT sketch matrices $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
5.3 Test results for Algs. 2.1 and 4.1
For Alg. 2.1 we output the relative error norms $\text{err}:=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$ for input matrices $M$, a target rank $r$, the $r$-truncation $M_{r}$ and $M(\rho)_{r}$, and the upper rank $\rho$, set to $2r,3r,4r$, $5r$ (cf. Remark 2.3).
[…]
Our tables display the mean of the relative error norms observed in these tests.
For Alg. 2.1 we also display the standard deviation.
For both algorithms we repeated the tests 100 times per test set with independent random choices of sketch matrices $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg. 2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg. 4.1 where TYPE stands for an input matrix type. We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT sketch matrices, which were more or less in the same range for both matrix types. All our tests with the synthetic input matrices from [28, Sec. 7.3.1] wit abridged SRHT sketch matrices have failed, that is, have output large error norms. In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random sketch matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
Alg. 2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT sketch matrices. This could have been predicted because the spectrum of the singular values $\sigma_{j}(M)$ of these matrices $M$ is flat for $j>20$ (decreasing roughly from $10^{-13}$ to between $10^{-15}$ and $10^{-16}$); in Table 5.1 we display the test results for the matrix Shaw and target rank $r=19$, which are about as favorable as for Gravity and Slow Decay.
In both cases of success and failure of both algorithms, our test results were in good accordance with our formal study. The relative error norms output by both Algs. 2.1 and 4.1 applied to $(r,\rho)$-matrices have consistently decreased and frequently to the optimal value 1 (up to the measurement errors) as we increased the upper rank $\rho$ for Alg. 2.1 and in the transition from the first to the second iteration of Alg. 4.1. In all tests except for High Noise, Slow Decay, and MinTemp matrices already err2 decreased below 1.1. The third iteration decreased the error norm much less or even slightly increased it.
For $(r,\rho)$-matrices the error norm bounds of Alg. 2.1 for $\rho=4r$ were consistently smaller than those of Alg. 4.1 in all three iterations. This was not the case just for the matrices with flat spectra such as the matrices Shaw for $r=20$, High Noise, and Exp Decay Slow.
Input Matrix Sketch Matrix $\rho$
$2r$ $3r$ $4r$ $5r$
Gravity Abr. SRHT 1.000 $\pm$ 1.036E-05 1.000 $\pm$ 7.815E-06 1.000 $\pm$ 7.888E-06 1.000 $\pm$ 6.683E-06
Gaussian 1.000 $\pm$ 7.495E-06 1.000 $\pm$ 5.211E-06 1.000 $\pm$ 7.214E-06 1.000 $\pm$ 5.432E-06
SLP Abr. SRHT 1.970 $\pm$ 4.083 1.000 $\pm$ 0.002 1.000 $\pm$ 1.018E-07 1.000 $\pm$ 1.362E-10
Gaussian 1.001 $\pm$ 0.0008160 1.000 $\pm$ 1.528E-07 1.000 $\pm$ 1.075E-10 1.000 $\pm$ 6.350E-14
Shaw, $r=19$ Abr. SRHT 1.000 $\pm$ 1.778E-06 1.000 $\pm$ 1.359E-06 1.000 $\pm$ 1.665E-06 1.000 $\pm$ 1.441E-06
Gaussian 1.000 $\pm$ 2.256E-06 1.000 $\pm$ 1.685E-06 1.000 $\pm$ 1.260E-06 1.000 $\pm$ 1.427E-06
Fast Decay Abr. SRHT 1.000 $\pm$ 3.086E-12 1.000 $\pm$ 3.590E-16 1.000 $\pm$ 2.470E-16 1.000 $\pm$ 2.617E-16
Gaussian 1.000 $\pm$ 6.651E-12 1.000 $\pm$ 5.117E-16 1.000 $\pm$ 4.552E-16 1.000 $\pm$ 4.590E-16
Slow Decay Abr. SRHT 1.000 $\pm$ 3.604E-05 1.000 $\pm$ 1.244E-06 1.000 $\pm$ 1.543E-07 1.000 $\pm$ 3.754E-08
Gaussian 1.000 $\pm$ 4.130E-05 1.000 $\pm$ 1.380E-06 1.000 $\pm$ 1.783E-07 1.000 $\pm$ 3.708E-08
Table 5.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, and synthetic matrices Fast Decay and Slow Decay, with fast and slowly decaying spectra.
Input Matrix Sketch Matrix 1st Itr. 2nd Itr. 3rd Itr.
Fast Decay Abr. SRHT 3.155000 1.0000 1.0000
Gaussian 3.1202 1.0000 1.0000
Slow Decay Abr. SRHT 5.0468 1.0003 1.0001
Gaussian 5.0755 1.0002 1.0001
Shaw for $r=20$ Abr. SRHT 2.8820E+01 1.0983 1.1225
Gaussian 1.8235E+01 1.1517 1.1189
Gravity Abr. SRHT 1.5762E+01 1.0000 1.0000
Gaussian 1.2917E+01 1.0000 1.0000
SLP Abr. SRHT 1.0931E+02 1.0014 1.0000
Gaussian 5.2205 1.0000 1.0000
Table 5.2: Test results for Alg. 4.1 on Fast Decay, Slow Decay, Shaw, Gravity, and SLP matrices.
Input Matrix $\rho$
$2r$ $3r$ $4r$ $5r$
Low Rank Low Noise 1.0416 $\pm$ 8.9977E-02 1.$0$$\pm$ 2.1197E-06 1.$0$$\pm$ 2.4039E-06 1.$0$$\pm$ 2.2834E-06
Low Rank Med Noise 1.4335 $\pm$ 1.7048E-01 1.0382 $\pm$ 3.1809E-02 1.0057 $\pm$ 1.4442E-03 1.0026 $\pm$ 5.8650E-04
Low Rank High Noise 5.6972 $\pm$ 8.6182E-01 4.8401 $\pm$ 4.3819E-01 4.0328 $\pm$ 2.3493E-01 3.7893 $\pm$ 2.1626E-01
Poly Decay Slow 2.0588 $\pm$ 1.8783E-01 1.6525 $\pm$ 2.0027E-01 1.3617 $\pm$ 8.6188E-02 1.2062 $\pm$ 8.6352E-02
Poly Decay Med 1.5384 $\pm$ 2.1907E-01 1.0315 $\pm$ 2.6585E-02 1.0028 $\pm$ 1.0190E-03 1.0009 $\pm$ 4.5652E-04
Poly Decay Fast 1.3133 $\pm$ 1.5431E-01 1.0001 $\pm$ 1.3887E-04 1.$0$$\pm$ 3.7489E-06 1.$0$$\pm$ 2.4495E-07
Exp Decay Slow 2.8587 $\pm$ 3.3389E-01 2.2772 $\pm$ 2.0481E-01 1.8244 $\pm$ 1.0970E-01 1.5721 $\pm$ 1.0528E-01
Exp Decay Med 1.5576 $\pm$ 1.2324E-01 1.0414 $\pm$ 4.9018E-02 1.0001 $\pm$ 9.0188E-05 1.$0$$\pm$ 3.7953E-07
Exp Decay Fast 1.3121 $\pm$ 1.4989E-01 1.$0$$\pm$ 6.4663E-11 1.$0$$\pm$ 3.6020E-16 1.$0$$\pm$ 3.0986E-16
Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [28, Sec. 7.3.1] using Gaussian sketch matrices.
Input Matrix 1st Itr. 2nd Itr. 3rd Itr.
Low Rank Low Noise 1.3940 1.0000 1.0000
Low Rank Med Noise 1.4752 1.0386 1.0375
Low Rank High Noise 1.4507 1.5135 1.5332
Poly Decay Slow 1.5920 1.4154 1.4073
Poly Decay Med 1.5569 1.0345 1.0306
Poly Decay Fast 1.3784 1.0001 1.0002
Exp Decay Slow 1.5202 1.4956 1.4750
Exp Decay Med 1.4946 1.0115 1.0164
Exp Decay Fast 1.4235 1.0000 1.0000
Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [28, Sec. 7.3.1] using Gaussian sketch matrices.
Input Sketch Matrix $\rho$
Matrix Type $2r$ $3r$ $4r$ $5r$
StreamVel Abr. SRHT 1.5840 $\pm$ 1.2583 1.4736 $\pm$ 1.4538 1.0022 $\pm$ 2.6478E-03 1.0002 $\pm$ 2.7733E-04
Gaussian 1.0932 $\pm$ 5.8887E-02 1.0064 $\pm$ 2.9809E-03 1.0013 $\pm$ 1.4141E-03 1.0003 $\pm$ 3.1762E-04
MinTemp Abr. SRHT 2.4575 $\pm$ 2.5748E-01 1.6980 $\pm$ 2.3998E-01 1.2919 $\pm$ 1.3606E-01 1.1181 $\pm$ 9.7415E-02
Gaussian 2.4393 $\pm$ 3.0823E-01 1.7216 $\pm$ 2.2510E-01 1.3793 $\pm$ 7.6629E-02 1.1238 $\pm$ 8.3331E-02
MaxCut Abr. SRHT 1.0024 $\pm$ 1.3630E-03 1.0005 $\pm$ 1.6875E-04 1.0002 $\pm$ 6.8092E-05 1.0001 $\pm$ 2.9040E-05
Gaussian 1.0026 $\pm$ 1.3856E-03 1.0005 $\pm$ 2.1129E-04 1.0002 $\pm$ 6.6260E-05 1.0001 $\pm$ 2.4207E-05
Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [28, Sec. 7.3.2]
Input Matrix Sketch Matrix Type 1st Itr. 2nd Itr. 3rd Itr.
StreamVel Abr. SRHT 2.9700 1.0739 1.0300
Gaussian 2.8724 1.0397 1.0177
MinTemp Abr. SRHT 2.3340 1.5961 1.5079
Gaussian 2.3120 1.6410 1.4950
MaxCut Abr. SRHT 3.6200E+01 1.0312 1.0182
Gaussian 2.6202E+01 1.0302 1.0176
Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [28, Sec. 7.3.2].
6 Conclusions
There are various natural extensions of our study, which can be practically competitive. E.g., one can devise and test combinations of Alg. 2.1 and of the TYUC and GN algorithms with various sparse sketch matrices such as Ultra-Sparse Rademacher sketch matrices [27, Sec. 3.9], sparsified random Givens rotations by Rokhlin and Tygert (see [10, Remark 4.6]), or even further sparsified count sketch matrices in Clarkson and Woodruff [2, 3].444[20, 21, 22] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes. Instead of random sampling LRA with sparse sketch matrices one can apply Cross-Approximation algorithm for LRA (see [17] and the references therein).
Our comparative study has showed no significant benefits of applying iterative refinement to LRA versus standalone Alg. 2.1, but one can further explore this comparison.
Appendix
Appendix A Small families of matrices that are hard for fast LRA
Any algorithm that does not access all entries of an input matrix fails to compute a close LRA of the following small families of matrices.
Example A.1.
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any algorithm that does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Such n algorithm outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. We arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Likewise, we can verify that any randomized LRA algorithms that does not access ail entries of an input matrix fails on those matrix families with error probability not close to 0.
Remark A.1.
How representative are the above matrix families of hard inputs? The matrices of these families represent data singularities and are not relevant to a large class of input matrices representing regular processes or, say, smooth surfaces. Furthermore, Example A.1 does not apply to various important special classes of input matrices such as symmetric or diagonally dominant.
Appendix B Computation of $r$-top SVD of a matrix product
For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [18, 24]) based on computation of the SVDs of $A$ and $B$.
Algorithm B.1.
[Computation of an $r$-top SVD of a matrix product.]
Input:
Four integers $r$, $k$, $m$, and $n$ such that $0<r\leq k\leq\min\{m,n\}$ and two matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$.
Output:
Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{C}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
Computations:
  1. 1.
    Compute SVDs $A=U_{A}\Sigma_{A}V^{*}_{A}$ and $B=U_{B}\Sigma_{B}V^{*}_{B}$ where $U_{B}\in\mathbb{C}^{m\times k}$, $V_{B}^{*}\in\mathbb{C}^{k\times n}$, and $\Sigma_{A},V^{*}_{A},U_{B},\Sigma_{B}\in\mathbb{C}^{k\times k}$.
  2. 2.
    Compute $k\times k$ matrices $W=\Sigma_{A}V^{*}_{A}U_{B}\Sigma_{B}$, $U_{W}$, $\Sigma_{W}$, and $V^{*}_{W}$ such that $W=U_{W}\Sigma_{W}V_{W}^{*}$ is SVD, $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$, and $\Sigma=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ and $\Sigma^{\prime}=\operatorname{diag}(\sigma_{j})_{j=r+1}^{k}$ are the matrices of the $r$ top (largest) and the $k-r$ trailing (smallest) singular values of the matrix $W$, respectively. Output the matrix $\Sigma$.
  3. 3.
    Compute and output the matrices $U$ and $V$ made up of the first $r$ columns of the matrices $U_{A}U_{W}$ and $V_{B}V_{W}$, respectively.
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [6, Figure 8.6.1]), running particularly fast where $k\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
Appendix C Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in our tests for Algs. 2.1 and 4.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [14, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (C.1)
For any fixed pair of $d$ and $i$, each of the matrices $H_{d,i}$ is orthogonal up to scaling and has $2^{d}$ nonzero entries in every row and column; we can compute the product $MH$ for an $n\times k$ submatrix $H$ of $H_{d,d}$ by using less than $km2^{d}$ additions and subtractions.
Now define the $d$-Abridged Scaled and Permuted Hadamard matrices, $PDH_{d,d}$, where $P$ is a random sampling matrix and $D$ is the matrix of random integer diagonal scaling. Each random permutation or scaling contributes up to $n$ random parameters. We can involve more random parameters by applying random permutation and scaling also to some or all intermediate matrices $H_{d,i}$ for $i=0,1,\dots,d$.
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [10, Sec. 11]).
Acknowledgement: Jianlin Xia helped place our study into proper historical context.